the sum of ages of a father and son is 70.in ten years time,the father will be twice as old as his son

Answer by Maths68(1474)



Let
Present age of father = f
Present age of son = s

The sum of ages of a father and son is 70.
f+s=70
f=70-s...........(1)

In ten years time
Father will be = f+10
Son will be = s+10
Then
The father will be twice as old as his son
f+10=2(s+10)
f+10=2s+20
f=2s+20-10
f=2s+10.................(2)
Put the value of f from (1) to (2)
70-s=2s+10
-s-2s=10-70
-3s=-60
-3s/-3=-60/-3
s=20
Put the value of s in (1)
f=70-s...........(1)
f=70-20
f=50
Present age of father = f = 50
Present age of son = s = 20